Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
OPERATION-2025
OPERATION-2024
OPERATION-2023
OPERATION-2022
OPERATION-2021
OPERATION-2020
OPERATION-2019
SASE1-2018
SASE1-2017
Experiment data
OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn
Not logged in
List
|
Message ID:
111
Entry time:
26 May 2024, 00:23
Author:
Romain Letrun
Group:
SPB
Subject:
Shift summary
X-ray delivery
6.0 keV, up to 202 pulses at 564 kHz and 101 at 282 kHz, 0.8-1.6 mJ
Low pulse energy after beam handover. Improved around 10:00.
~1.5 hour downtime due to issue with A5.
Several timing jumps of up to 1.5 ps visible on BAM and confirmed by X-ray/optical laser timing measurement at the instrument.
Optical laser delivery
770 nm, up to 202 pulses at 564 kHz and 101 at 282 kHz.
Stable delivery.
Achievements/Observations
.
Collected time-resolved diffraction from gold and gold-platinum nanorods at multiple laser fluences and over long and short delay scans.
Issues
AGIPD interlocked overnight and again during the morning due to temperature.
Had to perform timing measurement and correction multiple times throughout the day due to timing jumps.
ELOG V3.1.4-7c3fd00